# Comparison results for Markov tree distributions

### v1

@@ 5.2 A special property of star structures
$(V|U)\leq_{S}(V^{\prime}|U^{\prime})\quad\colon\Longleftrightarrow\quad F_{V|U=q_{U}(\cdot)}(v)\prec_{S}F_{V^{\prime}|U^{\prime}=q_{U^{\prime}}(\cdot)}(v)\quad\text{for all }v\in[0,1]\,,$ (37)
By definition, the Schur order for conditional distributions is invariant under rearrangements of the conditional variable and compares the variability of conditional distribution functions in the conditional variables in the sense of the Schur order for functions. Minimal elements characterize independence and maximal elements characterize complete directed dependence, see [4, Theorem 3.5]. Further, the Schur order for conditional distributions has the property that, for $U\stackrel{{\scriptstyle\mathrm{d}}}{{=}}U^{\prime}$ and $V^{\prime}\uparrow_{st}U^{\prime}\,,$
@@
i.e., less variability of the conditional distribution function $u\mapsto F_{V|U=u}(v)$ in the conditioning variable than $u\mapsto F_{V^{\prime}|U^{\prime}=u}(v)$ in sense of (36) for all $v$ implies that $(U,V)$ is smaller or equal in the supermodular order than $(U^{\prime},V^{\prime})$ whenever $V^{\prime}$ is stochastically increasing in $U^{\prime}\,,$ see [6, Proposition 3.17], cf. [4, Proposition 3.4]. Note that in (38), there is no positive dependence assumption on $(U,V)\,.$ If additionally $V\uparrow_{st}U\,,$ then also the reverse direction in (38) holds true, see [4, Proposition 3.4], cf. [6, 107]. In this case, the Schur order is equivalent to the supermodular order and we are back in the setting of Lemma 1.2.
The following result is a version of [8, Corollary 4(i)] and extends (38) to a vector of conditionally independent random variables. It states that a strengthening of the supermodular ordering condition to the Schur order allows to skip the positive dependence assumption (i) of Lemma 1.2 on $(X_{i},X_{0})\,.$
@@
Then it follows that $X\leq_{sm}Y\,.$ In particular, $Y$ is positive supermodular dependent.
In the following remark, we discuss the special properties of star structures which allow a more general supermodular comparison result based on the Schur order for conditional distributions.
@@
    Since $(V|U)\leq_{S}(V^{\prime}|U^{\prime})$ is equivalent to $(U,V)\leq_{sm}(U^{\prime},V^{\prime})$ whenever $U\stackrel{{\scriptstyle\mathrm{d}}}{{=}}U^{\prime}\,,$ $V\uparrow_{st}U$ and $V^{\prime}\uparrow_{st}U^{\prime}\,,$ Proposition 5.4 extends Lemma 1.2 to a large class of conditionally independent distributions that allow a non-positive dependence structure of $X\,.$ Surprisingly, as we show in Example A.6, Proposition 5.4 cannot be extended from star structures to Markov tree distributions: To be precise, let $T=(N,E)$ for $N=\{0,1,2,3\}$ and $E=\{(0,1),(1,2),(2,3)\}$ be a chain of $4$ nodes. Then, we can construct Markov tree distributed random variables $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ such that $(X_{i},X_{i+1})\leq_{sm}(Y_{i},Y_{i+1})$ and $(Y_{i},Y_{i+1})$ is $\mathrm{\text{MTP}_{2}}$ for all $i\in\{0,1,2\}$. Further, it holds that $(X_{i}|X_{j})\leq_{S}(Y_{i}|Y_{j})$ and $(X_{j}|X_{i})\leq_{S}(Y_{j}|Y_{i})$ for all $(i,j)\in E\,,$ as well as $X\not\leq_{lo}Y\,.$ Hence, less variability of the bivariate specifications $\{(X_{i},X_{j})\}_{(i,j)\in E}$ than $\{(Y_{i},Y_{j})\}_{(i,j)\in E}$ with respect to the Schur order as in (37) does not imply that the entire vector $X=(X_{n})_{n\in N}$ is smaller than $Y=(Y_{n})_{n\in N}$ with respect to the lower orthant order and thus neither with respect to the supermodular order, even if the specifications $(Y_{i},Y_{j})_{(i,j)\in E}$ satisfy the strong positive dependence concept $\mathrm{\text{MTP}_{2}}\,,$ see also Figure 8 d). Hence, Proposition 5.4 on the rearrangement-based Schur order is not extendable to Markov processes.

### v3

@@ 4 A special property of star structures
$(V|U)\leq_{S}(V^{\prime}|U^{\prime})\penalty 10000\ \penalty 10000\ \colon\Longleftrightarrow\penalty 10000\ \penalty 10000\ F_{V|U=q_{U}(\cdot)}(v)\prec_{S}F_{V^{\prime}|U^{\prime}=q_{U^{\prime}}(\cdot)}(v)\penalty 10000\ \penalty 10000\ \text{for all }v\in[0,1].$ (10)
By definition, the Schur order for conditional distributions is invariant under rearrangements of the conditioning variable. It compares the variability of conditional distribution functions in the conditioning variables in the sense of the Schur order for functions. Minimal elements characterize independence and maximal elements characterize perfect directed dependence; see [3, Theorem 3.5]. Further, the Schur order for conditional distributions has the property that, for $U\stackrel{{\scriptstyle\mathrm{d}}}{{=}}U^{\prime}$ and $V^{\prime}\uparrow_{st}U^{\prime},$
@@
i.e., less variability of the conditional distribution function $u\mapsto F_{V|U=u}(v)$ in the conditioning variable than $u\mapsto F_{V^{\prime}|U^{\prime}=u}(v)$ in sense of (9) for all $v$ implies that $(U,V)$ is smaller or equal in the supermodular order than $(U^{\prime},V^{\prime})$, provided $V^{\prime}$ is stochastically increasing in $U^{\prime};$ see [5, Proposition 3.17]. Note that in (11), there is no positive dependence assumption on $(U,V).$ If, additionally, $V\uparrow_{st}U,$ then the reverse direction in (11) also holds true; see [3, Proposition 3.4]. In this case, the Schur order is equivalent to the supermodular order, bringing us back to the setting of Lemma 1.2.
The following result is a version of [7, Corollary 4(i)] and extends (11) to a vector of conditionally independent random variables. It states that strengthening the supermodular ordering condition (iii) to the Schur order allows dropping the positive dependence assumption (i) of Lemma 1.2 on $(X_{0},X_{i}),$ $i\in\{1,\ldots,d\}.$
@@
Then it follows that $X\leq_{sm}Y.$ In particular, $Y$ is PSMD.
In the following remark, we discuss the distinctive property of star structures that enables the above generalization of Theorem 1.3 to the Schur order.
@@
    Since $(V|U)\leq_{S}(V^{\prime}|U^{\prime})$ is equivalent to $(U,V)\leq_{sm}(U^{\prime},V^{\prime})$ whenever $U\stackrel{{\scriptstyle\mathrm{d}}}{{=}}U^{\prime},$ $V\uparrow_{st}U$ and $V^{\prime}\uparrow_{st}U^{\prime},$ Proposition 4.2 extends Lemma 1.2 to a broader class of conditionally independent distributions that allow a non-positive dependence structure for $X.$ Surprisingly, as we show in Example 6.6, Proposition 4.2 cannot be extended from star structures to Markov tree distributions: To be precise, let $T=(N,E)$ for $N=\{0,1,2,3\}$ and $E=\{(0,1),(1,2),(2,3)\}$ be a chain of $4$ nodes. Then, we can construct Markov tree distributed random variables $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ such that
    • $(X_{i},X_{i+1})\leq_{sm}(Y_{i},Y_{i+1})$ for all $i\in\{0,1,2\}$,
    • $(Y_{i},Y_{i+1})$ being $\mathrm{\text{MTP}_{2}}$ for all $i\in\{0,1,2\}$, and
    • $(X_{j}|X_{i})\leq_{S}(Y_{j}|Y_{i})$ for all $(i,j)\in E,$
    but $X\not\leq_{lo}Y.$ Hence, we obtain that less variability of the bivariate specifications $\{(X_{i},X_{j})\}_{(i,j)\in E}$ compared to $\{(Y_{i},Y_{j})\}_{(i,j)\in E}$ with respect to the Schur order, as in (10), does not imply that the entire vector $X=(X_{n})_{n\in N}$ is smaller than $Y=(Y_{n})_{n\in N}$ in the lower orthant order and thus neither in the supermodular order—even if the specifications $(Y_{i},Y_{j})_{(i,j)\in E}$ satisfy the strong positive dependence concept $\mathrm{\text{MTP}_{2}};$ see also Figure 8 d). Consequently, Proposition 4.2 on the rearrangement-based Schur order has no direct extension to Markov tree distributions.